# Effect Identification and Unit Categorization in the Multi-Score Regression Discontinuity Design with Application to LED Manufacturing

### v1

@@ 5.1 Set-Up
For this, we generate semi-synthetic data following Algorithm 1. The python implementation of this process as well as all estimators is publicly available. The process is calibrated to match real data characteristics. We draw n=10,000n=10{,}000 observations and repeat each experiment $r=250$ times.
We evaluate the cut-offs $c_{D}$ and $c_{Y}$ separately, estimating the complier effect of $G\in\{I_{Y},\,I_{D}\}$ as identified in Theorem 1, the subset complier effect of $G$ given its counterpart and intent-to-treat estimates with and without the subset conditioning. The complier effects are estimated under a fuzzy design; the intent-to-treat effects under a sharp design.
The oracle values are estimated using local linear kernel regression on the differences in true potential outcomes. The covariates consist of statistics describing the quality of individual items. We use different estimators including covariate adjusted estimators estimating the complier and subset complier effect
Table 1 provides an overview of the estimators we use.
@@ 5.2 Estimators
The basic RDD estimator runs separated local linear regressions on each side of the cutoff:
@@
where the $w_{i}(h)$ are local linear regression weights that depend on the data through the realizations of the running variable only and $h>0$ is a bandwidth.
Under standard conditions, which include that the running variable is continuously distributed, and that the bandwidth $h$ tends to zero at an appropriate rate, the estimator $\hat{\tau}_{\text{base}}(h)$ is approximately normally distributed in large samples, with bias of order $h^{2}$ and variance of order $(nh)^{-1}$.
The conventionally adjusted estimator includes covariates linearly in the local regressions. We also use modern RDD estimators with flexible covariate adjustment based on potentially nonlinear adjustment functions $\eta$. The estimator takes the following form:
$\widehat{\tau}_{\text{RDFlex}}(h;\eta)=\sum w_{i}(h)M_{i}(\eta),\quad M_{i}(\eta)=Y_{i}-\eta(Z_{i}).$ (10)
$\eta$ is the influence of $Z$ on the outcome $Y$ and is estimated by ML methods.
@@ 5.3 Cautious Operator
Figure 7 depicts a side-by-side the estimators for the effect at $c_{D}$ and $c_{Y}$ for the case of the cautious operator. The intent-to-treat oracles are closer to zero than the complier effects due to the inclusion of individuals who are nevertakers with respect to each cutoff rule. As visible in 7(a), for $I_{D}$ we estimate an overall negative effect, although it is not significant at $95\%$-level. The subset effect for the fuzzy case exhibits a smaller bias as the percentage of nevertakers in the estimation sample is smaller. The estimated effect remains unchanged as only nevertakers but no compliers were removed. For the intent-to-treat estimator, a higher share of compliers in the subsample increases the estimated effect of treatment rule $G$. The subset estimators have a comparable estimated variance (see Table 2), with the coverage overall appearing slightly more credible. Generally, the covariate adjustment reduces the standard error in the estimation, especially for the sharp estimators. The estimates for $I_{Y}$ are small and positive. The fuzzy estimator on the full data has a high standard error, which increases further with ML adjustment. This may be due to a small jump in treatment probability in the full data, destabilizing the ML estimate. The subset estimator along this axis removes much more observations within the bandwidth of the estimation thus decreasing the variance. Additional results can be found in Appendix B.
@@
Figure 7: Comparison of median estimates with median 95% confidence interval at $c_{D}$ and $c_{Y}$ on the simulated data with “cautious” operator assumption. Different colors depict complier effect (fuzzy) as well as intent-to-treat effect (sharp). The subset estimator excludes nevertaker according to Theorem 2.
@@
It is worth mentioning some results of two-dimensional estimators as suggested in previous literature. We compare the “binding-the-score” approach, that normalizes and combines $X_{D}$ and $X_{Y}$ into a singular score component and the “Euclidean distance” approach, that computes the shortest distance of each observation to a predetermined point (in our case a point on $c_{D}$.
Figure 8: Median estimates with median 95% confidence interval Comparison of median estimates with median 95% confidence interval Estimators that use two dimensions at once.
The results in Figure 8 show oracles very close to zero. This indicates that when collapsing both dimensions of the decision rule into one it is no longer possible to have individual insights into the rules. Interestingly, the estimator that uses the euclidean distance to a boundary point exhibits a significant bias across all methods. In our specific setting, there is no straightforward interpretation of the RDD effect estimate at a two-dimensional threshold, which contrasts with previous work where score components are more directly comparable, such as test scores or geographical distances.
@@ 5.4 Acknowledging Operator
We compare the results from the previous section to the “acknowledging” operator, who accepts the the decision of the system without using his additional domain knowledge. Thus, in this setting the data follows a sharp two-dimensional MRD. The comparison of the effects at $c_{Y}$ and $c_{D}$ can be seen in Figure 9
@@
Figure 9: Comparison of median estimates with median 95% confidence interval at $c_{D}$ and $c_{Y}$ on the simulated data with “acknowledging” operator assumption. Different colors depict complier effect (fuzzy) as well as intent-to-treat effect (sharp). The subset estimator excludes nevertaker according to Theorem 2
In this scenario, the effects are smaller for $c_{D}$, as lots that the cautious operator did not rework due to his domain knowledge about possible degradation, are reworked. This leads to an overall larger negative effect and suggests moving the threshold even further than in the scenario above. The sharp subset estimator estimates the identical effect as the fuzzy estimator on the full data, as subsetting the data according to $I_{Y}$ leads to a sharp design with all individuals complying to threshold movements. The subset estimators estimate the effect overall with a slightly smaller bias and standard error. The sharp estimator on the full sample, which corresponds to an intent-to-treat estimator, estimates an effect closer to zero, as it includes nevertakers.
At the cutoff $c_{Y}$, the change of behavior of the operator does not change the effect estimates much. This can be explained that less individuals in the neighborhood of the cutoff are effected by this.
@@ 6 Empirical Results
This section presents results from estimation on the real data with n=9,103n=9{,}103 observations from the production system. Additional to the covariates used above, we include the shopfloor workload. Our primary interest in the real system is evaluating $I_{D}$.
As shown in Table 3, we estimate a positive effect. The subset estimates have a lower variance, as we remove uncertainty due to nevertakers according to $I_{Y}$. In case of the intent-to-treat estimators, the addition of ML adjustment improves the variance of the estimation. For the fuzzy estimators, there is no clear improvement.
@@
The comparison in Figure 10 provides insight into the differing effect signs between semi-synthetic and real data. While the semi-synthetic process (left) appears to have a rework threshold closer than the optimal distance, the real data (right) suggests a threshold that is further away, resulting in a significant outcome jump at the current threshold.
@@
Figure 10: Comparison of local linear regressions around the cutoff for the semi-synthetic and real data. The opposite sign of the estimated effect can be explained by the calibration of the semi-synthetic process.
@@
We validate our results from the application with a pseudo-cutoff test (see Cattaneo et al. (2019)). We compare the estimated intent to treat effect of the real data at $c_{D}$ with a fictional lower and a higher cutoff.
In Table 4 it is visible that both pseudo cutoffs show effects close to zero and with large confidence band, while at the real cutoff the effect is almost significant on $95\%$ confidence level.
@@
Table 4: Coefficient and confidence interval of the sharp real data estimation at pseudo cutoffs concerning $c_{D}$.
@@ 7 Conclusion
With this paper we present progress towards better applicability of the regression discontinuity design in Economics and Management. The multi-score RDD framework combined with modern ML adjustment yields great opportunities to evaluate complex decision settings and policies. Particularly, in settings with strict decision boundaries in place, policy makers experience little to no overlap, limiting the possibility of “what happens if”-reasoning.
Our theoretical insights are based on three intuitive observations. First, the analysis of more general assignment rules requires lifting the concept of “binary compliance” to the higher-level concept of “compliance with an assignment mechanism”. That is, in theory all possible inputs of the mechanism need to be considered when assessing compliance. Secondly, and most evident by the first observation, the concept of compliance is only relative to the assessed rules and thus changes of the assignment mechanism can alter the perceived unit behavior. A fact that should be reflected in the theory as knowledge about parts of the assignment mechanism and/or the final decision-rule allows identifying noncompliance. Lastly, in multiple dimensions the complier effect at the cutoff should not depend on the direction from which it is approached, a property that can be exploited by future estimation methods. Formalizing these observations we generalized existing unit behavior types (complier, nevertaker, alwaystaker) for multi-dimensional cutoff-rules fully characterizing the possible behavior types with the introduction of indecisive items. We derived rules that show how the unit behavior changes when cutoff-rules are decomposed into sub-rules. Further we proved an identification result for the cutoff complier effect in MRD settings, and provided conditions for the stability of this result when subsets of alwaystaker and nevertaker are removed, giving rise to the sub-set complier effect at the cutoff.
@@
We validated our theory both with simulation studies using a semi-synthetic DGP and in a real-world application setting from manufacturing, showing that that subrule-specific estimation provides insights on how to improve existing decision-making. In particular the results confirm that removing identifiable nevertaker and/or alwaystaker allow for more precise effect estimation, under conditions provided by the theory. Combined with flexible ML-based adjustments variance can be further reduced without introducing additional bias. We advise the practitioner to estimate both the complier and the sub-set complier effect at the cutoff whenever nevertaker and/or alwaystaker are identifiable in application, and retrospectively validate the assumptions. Importantly, the empirical findings underline the asymmetry in the treatment effects across different cutoff dimensions, highlighting that effect of the whole assignment rule is in general not equal to the sub-rule effects, as is evident by the theory. Nevertheless, a nuanced evaluation of each score component, can provide valuable insights.
@@
We expanded the existing MRD theory providing means for the analysis of complex cutoff-rules, and their corresponding sub-rules. For this we used the language of well-defined unit behavior types. While we bridged the gap between econometric RDD methods and industrial process control, our results remain broadly applicable in complex decision settings. Future work could explore the relation of main- and sub-rules more closely, expand on the stability of unit types, or extend the multi-dimensional RDD setting to hierarchical decision structures across multiple stages. Additionally, developing estimators tailored to the identification theorems are promising future avenues.
@@ A.1 Unit categorization
See 1 Before proofing the above statement let us make some simple observations. Let
@@ A.2 Examples
Nevertaker, Complier, Alwaystaker
We now investigate potential setup for cutoff rules that admit complier, alwaystaker and nevertaker. In accordance with the empirical part let $T\coloneqq I_{1}\square I_{2}$ for $\square\in\{\lor,\land\}$ and $G=I_{k}$ for $k=1,2$ and further let $D$ be a cutoff rule. Thus, only the number of additional variables in $D$ limit the possible unit categories. That is, with one additional cutoff variable $I_{3}$ one could realize each tuple of the different item types (e.g. defier and nevertaker). A common assumption in causal literature is that there are no defying or indecisive units. Thus, to have a setting with each of the remaining three categories one would require one more variable $I_{4}$. This leaves a single choice open (nevertaker, alwaystaker or complier) for the remaining region in $\mathbb{R}^{4}$. As in case of the cautious operator we decide to opt for compliance. Thus, we arrive at
@@
That is $X_{4}$ could be ignored for estimating the complier effect of $G$ due to our the specific choice, whereas $X_{3}$ and $X_{2}$ are the switches required to identify $\operatorname{ComP}(G,D)$ completely. Again $X_{2}$ might be known from $T$, which helps to reduce the observational dataset accordingly.
@@ Acknowledgements
This work was funded by the Bavarian Joint Research Program (BayVFP) – Digitization (Funding reference: DIK0294/01). The research partners ams OSRAM and Economic AI kindly thank the the VDI/VDE-IT Munich for the organization and the Free State of Bavaria for the financial support.

### v2

@@ 5.1 Set-Up
We generate semi-synthetic data following Algorithm 1. A Python implementation of the DGP and all estimators is publicly available. The process is calibrated to match the characteristics of real production data. We draw n=10,000n=10{,}000 observations and repeat each experiment $r=250$ times.
We evaluate the cut-offs $c_{D}$ and $c_{Y}$ separately, estimating:
  • the complier effect of $G\in\{I_{Y},\,I_{D}\}$, as identified in Theorem 1;
  • the subset complier effect of $G$, conditional on its counterpart; and
  • intent-to-treat (ITT) effects, with and without the subset conditioning.
Complier effects are estimated under a fuzzy design, whereas ITT effects are estimated under a sharp design. Oracle values are obtained using local linear kernel regression on the differences in true potential outcomes. Covariates consist of statistics describing the quality of individual items.
We consider a variety of estimators, including covariate-adjusted estimators, for complier and subset complier effects. Table 1 provides an overview of the estimators used in our analysis.
@@ 5.2 Estimators
The basic RDD estimator runs separate local linear regressions on each side of the cutoff:
@@
where the $w_{i}(h)$ are local linear regression weights that depend on the data through the realizations of the running variable only, and $h>0$ is a bandwidth.
Under standard conditions (e.g., the running variable is continuously distributed and the bandwidth $h$ tends to zero at an appropriate rate), the estimator $\hat{\tau}_{\text{base}}(h)$ is approximately normally distributed in large samples, with bias of order $h^{2}$ and variance of order $(nh)^{-1}$.
A conventional extension is the covariate-adjusted estimator, which incorporates covariates linearly into the local regressions. We also use modern RDD estimators with flexible covariate adjustment based on potentially nonlinear adjustment functions $\eta$. This estimator takes the form:
$\widehat{\tau}_{\text{RDFlex}}(h;\eta)=\sum_{i=1}^{n}w_{i}(h)M_{i}(\eta),\quad M_{i}(\eta)=Y_{i}-\eta(Z_{i}),$ (10)
where $\eta$ denotes the influence of $Z$ on the outcome $Y$, estimated using machine learning methods.
@@ 5.3 Cautious Operator
Figure 7 presents side-by-side results for the estimators at $c_{D}$ and $c_{Y}$ in the case of the cautious operator. The intent-to-treat (ITT) oracles are closer to zero than the complier effects because they include individuals who are nevertakers with respect to each cutoff rule. As shown in Figure 7(a), for $I_{D}$ we estimate an overall negative effect, although it is not statistically significant at the 95% level.
The subset effect for the fuzzy case exhibits a smaller bias, since the proportion of nevertakers in the estimation sample is lower. The estimated effect remains unchanged because only nevertakers, but no compliers, were removed. For the ITT estimator, a higher proportion of compliers in the subsample increases the estimated effect of treatment rule $G$. The subset estimators have a comparable variance (see Table 2), with coverage appearing slightly more credible overall. In general, covariate adjustment reduces standard errors, especially for the sharp estimators.
The estimates for $I_{Y}$ are small and positive. The fuzzy estimator on the full data has a high standard error, which increases further with ML adjustment. This may be due to a small jump in treatment probability in the full data, destabilizing the ML estimate. In contrast, the subset estimator along this axis removes more observations within the estimation bandwidth, thereby reducing variance. Additional results can be found in Appendix B.
@@
Figure 7: Comparison of median estimates with median 95% confidence intervals at $c_{D}$ and $c_{Y}$ on the simulated data with the “cautious” operator assumption. Different colors depict complier effect (fuzzy) as well as intent-to-treat effect (sharp). The subset estimator excludes nevertakers according to Theorem 2.
@@
Table 2: Mean bias, standard error, coverage, and first-stage prediction quality of different estimators in estimation at $c_{D}$.
@@
We also consider two-dimensional estimators proposed in previous literature. We compare the “binding-the-score” approach, which normalizes and combines $X_{D}$ and $X_{Y}$ into a single score, and the “Euclidean distance” approach, which computes the shortest distance of each observation to a predetermined point (here, a point on $c_{D}$).
Figure 8: Median estimates with median 95% confidence intervals for estimators that use two dimensions at once.
The results in Figure 8 show oracles very close to zero, indicating that when collapsing both dimensions of the decision rule into one, it is no longer possible to obtain insights into the individual rules. Interestingly, the estimator based on the Euclidean distance exhibits a significant bias across all methods.
In our specific setting, there is no straightforward interpretation of the RDD effect estimate at a two-dimensional threshold, which contrasts with previous applications where score components are more directly comparable, such as test scores or geographical distances.
@@ 5.4 Acknowledging Operator
We now compare the results from the previous section to those for the “acknowledging” operator, who accepts the decision of the system without using additional domain knowledge. In this setting, the data follow a sharp two-dimensional MRD. The comparison of the effects at $c_{Y}$ and $c_{D}$ is shown in Figure 9.
@@
Figure 9: Comparison of median estimates with median 95% confidence intervals at $c_{D}$ and $c_{Y}$ on the simulated data with the “acknowledging” operator assumption. Different colors depict complier effect (fuzzy) as well as intent-to-treat effect (sharp). The subset estimator excludes nevertakers according to Theorem 2.
For $c_{D}$, the estimated effects are smaller than under the cautious operator. Lots that the cautious operator withheld from rework due to concerns about degradation are reworked in this case. This leads to a larger overall negative effect and suggests that the threshold should be moved even further than in the cautious operator case. The sharp subset estimator yields the same effect as the fuzzy estimator on the full data, because subsetting according to $I_{Y}$ produces a sharp design in which all individuals comply with threshold movements. Across estimators, the subset versions display slightly smaller bias and standard errors. In contrast, the sharp estimator on the full sample, which corresponds to an ITT estimator, produces an effect closer to zero because it includes nevertakers.
At the cutoff $c_{Y}$, the operator’s behavior has little impact on effect estimates. This can be explained by the fact that relatively few individuals in the neighborhood of the cutoff are affected.
@@ 6 Empirical Results
This section presents results from estimation on the real data, which consist of n=9,103n=9{,}103 observations from the production system. In addition to the covariates used above, we include shop-floor workload. Our primary interest in the real system is evaluating the rule $I_{D}$.
As shown in Table 3, we estimate a positive effect. The subset estimates display lower variance because uncertainty from nevertakers (according to $I_{Y}$) is removed. For the ITT estimators, ML adjustment improves variance, whereas for the fuzzy estimators no clear improvement is observed.
@@
The comparison in Figure 10 highlights differences between the semi-synthetic and real data. The semi-synthetic results (panel a) suggest that the rework cutoff is miscalibrated, leading to negative estimated effects, whereas the real data (panel b) show a positive jump at the current cutoff, consistent with a beneficial rework threshold.
@@
Figure 10: Comparison of local linear regressions around the cutoff for the semi-synthetic (a) and real (b) data. The opposite sign of the estimated effects can be explained by calibration of the semi-synthetic process.
@@
We validate the application results using a pseudo-cutoff test (see Cattaneo et al. (2019)). We compare the estimated ITT effect at the true cutoff $c_{D}$ with estimates at a hypothetical lower and higher cutoff.
As shown in Table 4, both pseudo cutoffs produce effects close to zero with wide confidence intervals, whereas the effect at the true cutoff is nearly significant at the 95% level.
@@
Table 4: Coefficient and confidence interval of the sharp real-data estimation at pseudo cutoffs for $c_{D}$.
@@ 7 Conclusion
This paper improves the applicability of regression discontinuity design (RDD) in operations and management. The multi-score RDD framework, combined with modern machine-learning-based adjustment, provides new opportunities to evaluate complex decision settings and policies. In settings with strict decision boundaries, policy-makers face little to no overlap in treatments near the cutoff, limiting the scope for “what if” reasoning.
Our theoretical insights rest on three observations. First, analyzing general assignment rules requires extending the notion of “binary compliance” to the broader concept of “compliance with an assignment mechanism.” In principle, all possible inputs to the mechanism must be considered when assessing compliance. Second, compliance is always defined relative to the rules under consideration. Changing the assignment mechanism can change the observed behavior of units. This relativity should be reflected in theory, as knowledge about specific parts of the mechanism or the final decision rule makes it possible to identify noncompliance. Third, in multiple dimensions, the complier effect at the cutoff should not depend on the direction from which the cutoff is approached, a property that future estimation methods can exploit.
Building on these observations, we generalized existing unit behavior types (compliers, nevertakers, alwaystakers) to multi-dimensional cutoff rules and introduced a new category of indecisive units. We derived rules describing how behavior types evolve when cutoff rules are decomposed into subrules. We also proved an identification result for the cutoff complier effect in MRD settings and established conditions under which this result remains valid when subsets of alwaystakers and nevertakers are removed, leading to the subset complier effect at the cutoff.
@@
We validated our theoretical results using simulation studies with a semi-synthetic DGP and a real-world application in manufacturing. These analyses show that subrule-specific estimation can improve decision-making. In particular, the results confirm that removing identifiable nevertakers and/or alwaystakers makes effect estimation more precise, under conditions specified by the theory. When combined with flexible ML-based adjustments, variance can be further reduced without introducing bias. We recommend that practitioners estimate both the complier and the subset complier effect at the cutoff whenever nevertakers or alwaystakers can be identified in practice, and retrospectively validate the required assumptions. Importantly, the empirical findings underline asymmetries in treatment effects across different cutoff dimensions: the effect of the overall assignment rule generally differs from the effects of its subrules. Nevertheless, careful evaluation of each score component can provide valuable insights.
@@
We expanded MRD theory by introducing tools to analyze complex cutoff rules and their corresponding subrules, expressed in terms of well-defined unit behavior types. Although our empirical study specifically bridges econometric RDD methods and industrial process control, the framework is broadly applicable to other complex decision settings, such as multi-criteria loan approval in financial institutions or public health resource allocation where eligibility depends on several risk factors. Future work could explore the relationship between main rules and subrules in greater depth, study the stability of unit types, or extend the multi-dimensional RDD to hierarchical decision structures across multiple stages. Developing estimators tailored to the identification theorems also represents a promising direction.
@@ A.1 Unit categorization
See 1 Before proving the above statement let us make some simple observations. Let
@@ A.2 Examples
Nevertakers, Compliers, Alwaystakers
We now investigate potential setup for cutoff rules that admit compliers, alwaystakers and nevertakers. In accordance with the empirical part let $T\coloneqq I_{1}\square I_{2}$ for $\square\in\{\lor,\land\}$ and $G=I_{k}$ for $k=1,2$ and further let $D$ be a cutoff rule. Thus, only the number of additional variables in $D$ limit the possible unit categories. That is, with one additional cutoff variable $I_{3}$ one could realize each tuple of the different item types (e.g. defier and nevertakers). A common assumption in causal literature is that there are no defying or indecisive units. Thus, to have a setting with each of the remaining three categories one would require one more variable $I_{4}$. This leaves a single choice open (nevertakers, alwaystakers or compliers) for the remaining region in $\mathbb{R}^{4}$. As in case of the cautious operator we decide to opt for compliance. Thus, we arrive at
@@
That is $X_{4}$ could be ignored for estimating the complier effect of $G$ due to our specific choice, whereas $X_{3}$ and $X_{2}$ are the switches required to identify $\operatorname{ComP}(G,D)$ completely. Again $X_{2}$ might be known from $T$, which helps to reduce the observational dataset accordingly.
@@ Acknowledgements
This work was funded by the Bavarian Joint Research Program (BayVFP) – Digitization (Funding reference: DIK0294/01) and the German Research Foundation (DFG) (Grant no 448795504). The research partners ams-OSRAM and Economic AI kindly thank the the VDI/VDE-IT Munich for the organization and the Free State of Bavaria as well as the Federal Republic of Germany, for the financial support.